# ppa_t_p.bib

### v1

@@ Abstract
With the increasing use of online services, the protection of the privacy of users becomes more and more important. This is particularly critical as authentication and authorization as realized on the Internet nowadays, typically relies on centralized identity management solutions. Although those are very convenient from a user’s perspective, they are quite intrusive from a privacy perspective and are currently far from implementing the concept of data minimization. Fortunately, cryptography offers exciting primitives such as zero-knowledge proofs and advanced signature schemes to realize various forms of so-called anonymous credentials. Such primitives allow to realize online authentication and authorization with a high level of built-in privacy protection (what we call privacy-preserving authentication). Though these primitives have already been researched for various decades and are well understood in the research community, unfortunately, they lack widespread adoption. In this paper, we look at the problems, what cryptography can do, some deployment examples, and barriers to widespread adoption. Latter using the example of the EU Digital Identity Wallet (EUDIW) and the recent discussion and feedback from cryptography experts around this topic. We also briefly comment on the transition to post-quantum cryptography.
@@ 1 Introduction
Authentication and authorization are two closely coupled and important tasks that are frequently required when users want to access resources in the digital realm, e.g., services on the Internet. Loosely speaking, authentication is the task of confirming whether a user is really who they pretend to be, e.g., by demonstrating the knowledge of some secret such as a password or a secret signing key associated to the user. Today, one increasingly uses a combination of more than one factor (multi-factor authentication), like the knowledge of a password and the possession of Subscriber Identity Module (SIM) card. Authorization is the task of giving users permissions to access certain resources, i.e., deciding whether a certain authenticated users has sufficient permissions to perform some action. Henceforth, we will just talk about authentication and note that typically authorization can be performed based on information associated to the identity after authentication and without any further interaction with the user. A typical example is the use of attributes associated to users, e.g., any user with attribute role=Student might be granted access to a certain service.
@@
A convenient way to think about a digital identity is that of a set of attributes that describe a person. Now, in a typical scenario a person’s identity comprises several partial identities (cf. [anonterminology]), where each of them represents a person in a specific context or role and can be thought of as a certain subset of attributes. The complete identity of a person is represented by the union of all attributes of all the partial identities of this person. Since some of these partial identities include sensitive information, from a privacy perspective it is desirable to have full control over which attributes are revealed in a certain context. This is particularly relevant in the digital realm where it is easy to collect such information, connect it to other information sources and use it to build up ever increasing profiles of individuals.
@@
Limiting the amount of information that is provided as much as possible is known as data minimization. It is a concept of growing importance in a globally connected digitized world and a fundamental principle behind many data protection regulations such as the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States. This helps to preserve privacy and can help to ensure as much unlinkability as possible, making it harder to easily connect all the partial identities of a person. One issue that is important to stress is that for many services on the Internet, concretely identifying information such as name or DateOfBirth of a user might not be required in the process of authentication, e.g., an individual can demonstrate that they are old enough to consume a certain service. What however can be required in various cases such as governmental applications, is that actions of the same partial identity of a person can be linked over different sessions. This can be achieved via so called pseudonyms, which represent identifiers other than the person’s real name (or other uniquely identifying information) that are independent of the person and provide a sufficient degree of anonymity.
@@
After having introduced some basic concepts, our goal is to look into how authentication on the Internet typically looks today and how it could look when putting privacy as the main design goal. We will discuss issues with these existing mechanism and present a concept from the cryptographic literature called anonymous credentials. This is an important tool to realize privacy-preserving authentication first envisioned by Chaum [DBLP:journals/cacm/Chaum85] and later realized by Camenisch and Lysyanskaya [DBLP:conf/eurocrypt/CamenischL01] as well as Brands [BrandsBook] in the early 2000s. After around 25 years, there is a large body of research on anonymous credentials and related primitives (cf. [DBLP:conf/secsr/KakviMPQ23, cryptoeprint:2023/1039]). Unfortunately, although this concept is well known in the research community, and over the recent years increasingly attracting practitioners, it still lacks a widespread deployment. Besides providing a high-level discussion of the theory of anonymous credentials, this paper will also look at the practice of anonymous authentication and potential issues that are hindering a widespread adoption. As a supporting use-case, we take the European Digital Identity Wallet (EUDWI), its architectural reference framework (ARF) and the recent feedback given by cryptographic experts [cryptoeudi], which highlights obstacles on the way to deploying such a “new” technology into an existing (legacy) infrastructure.
@@
Outline of this paper. In Section 2 we review traditional authentication on the Internet. Then, in Section 3 we introduce relevant privacy properties and discuss privacy of traditional authentication approaches. In Section 4 we introduce the concept of anonymous credentials from a conceptual (theoretical) perspective. Then, in Section 5 we discuss anonymous credentials from a practical perspective and in particular implementations, standardization efforts and real-world use-cases. In Section 6 we take a look at the European Digital Identity Wallet (EUDWI), its planned realization and related privacy issues. Then in Section 7 we look at barriers and issues to be considered when aiming at deploying anonymous credentials such as integrating them into the EUDWI. Finally, in Section 8 we conclude this paper and give an outlook on future aspects related to deploying anonymous credentials.
@@ 2 Traditional Authentication and Identity-Management
The very traditional way of authenticating users on the Internet is that every service realizes its own authentication mechanism, meaning that a user establishes a partial identity with every single service (typically pseudonyms or more likely email addresses). These partial identities are though typically not distinct, as users tend to re-use their email addresses either as explicit pseudonym or as an attribute given to the service provider. Latter is usually required for functionality reasons, i.e., account recovery.
@@ 2.1 Password-based Authentication
The most common traditional authentication mechanism is via user-chosen passwords associated to user accounts. Since this is not very convenient, it is well known that users tend to choose too weak passwords [DBLP:conf/sp/Bonneau12] as well as reuse them among different services [DBLP:conf/soups/WashRBW16]. While there are password-management tools that help to overcome these issues, research shows that the use of such tools significantly lacks behind [DBLP:conf/soups/ZhangPBC19] and due to various reasons does not see widespread use. Despite all these issues, due to its simplicity and easy deployment it is still the most common authentication mechanism.
@@ 2.2 Signature-based Authentication
From a security perspective more desirable and increasingly pushed, is the use of a particular two-factor (2F) authentication relying on strong cryptography. A popular choice is the use of the FIDO2222https://fidoalliance.org/fido2/ de-facto standard from the Fast IDentity Online (FIDO) alliance. Here one either uses a dedicated hardware token or a pure software component (typically called Passkey [^3] https://fidoalliance.org/passkeys/) as the second factor. Latter can be easily used with any device like a smartphone and thus increases the usability.
@@
The technical protocol details are not important here, but the relevant part is that it uses public-key cryptography and in particular digital signatures. We recall that in context of a digital signature scheme every user can generate a pair $(\mathsf{sk},\mathsf{pk})$ consisting of a secret (signing) key $\mathsf{sk}$ and a public (verification) key $\mathsf{pk}$, where latter is made public. A user knowing $\mathsf{sk}$ can produce a signature $\sigma\leftarrow\mathsf{Sign}(\mathsf{sk},m)$ for any message $m$, and using $\mathsf{pk}$ and given $(m,\sigma)$ anyone can check the validity of the signature using $\{0,1\}\leftarrow\mathsf{Verify}(\mathsf{pk},m,\sigma)$, where $1$ indicates a valid signature. A valid signature gives the guarantees that $m$ has really been signed by the holder of the secret key corresponding to $\mathsf{pk}$ and has not been altered in any way. Security requires that a signature scheme is unforgeable. This means that when only having $\mathsf{pk}$ and access to signatures for arbitrarily chosen messages, it is not possible to produce a valid signature $\sigma$ for some message $m^{\prime}$ that has not been explicitly signed by the signer, i.e., asked a signature for.
@@
In conventional web authentication, $\mathsf{pk}$ is typically just stored together with the identity and thereby establishing the binding between the individual and the key. More generally, this binding between $\mathsf{pk}$ and an identity is achieved via explicit certification. The most common way is to rely on a public-key infrastructure (PKI). This means that there is some trusted certification authority (CA) that signs $\mathsf{pk}$ together with a set of attributes of the user, resulting in a certificate that can be verified by anyone trusting the public key of the CA. Latter is typically used for signature-based authentication when using an official governmental electronic identity (eID), as largely deployed within the European Union, where some governmental body acts as a CA.
2.3 Single-Sign-On
@@
A widespread concept used nowadays is that of single sign-on (SSO), which makes account management simpler. Here a user is required to authenticate only at a single centralized entity (typically a large technology company such as Google, Microsoft or Meta) acting as an identity-provider (IdP) for the user. Then one employs an authentication and authorization layer such as OpenID Connect (OIDC), which enables authentication at third party services (so called relying parties or RPs) via the IdP. We illustrate the process in Figure 1.
@@
Figure 1: Abstract concept of single-sign-on.
Essentially, whenever a user wants to authenticate at an RP (1), the user is re-directed to it’s IdP (2) where the user actually authenticates using e.g., password- or signature-based authentication (3). After a successful authentication, the IdP generates a token (an ID Token in case of OIDC), signed by the IdP (4), which the user can hand over to the RP (5) as a proof that the user has authenticated to the IdP.
@@ 3 Privacy Properties and Traditional Approaches
Now having discussed ways of authenticating users on the Internet, we want to discuss privacy properties that might be desirable from such mechanisms.
@@ 3.1 Privacy Properties
We will briefly summarize the privacy properties that are most important in context of authentication at an intuitive level. We will not take into account other important security properties of authentication mechanisms which are not relevant to privacy, e.g., that it is hard to impersonate other users. Also in our further consideration we will not consider privacy on the network layer, i.e., identifiers such as IP addresses or other meta data on the network layer that might be used to identify users. There are overlay networks [DBLP:conf/sp/SyversonGR97, DBLP:conf/uss/DingledineMS04] and in particular Tor [^4] https://www.torproject.org to achieve anonymity. They can be used to hide such information and obtain anonymity on the network layer in practice. For simplicity we just assume that using such techniques, anonymity is realized in an ideal way. But we stress that in practice for such networks to be effective there are many issues that need to be considered [DBLP:conf/diau/Raymond00]. Now we list the properties:
@@
Technically, unlinkability means that different actions (or transactions) of a user cannot be linked together. We will use the terms IdP, even though it might be a certificate issuer only, and RPs as a generic term for service providers. Since in such a setting there are different parties involved, we have to consider unlinkability in settings where different of these parties collaborate in trying to break the unlinkability. Below, we will use the terminology in [cryptoeudi]:
@@
If a user authenticates to different RPs, those RPs cannot determine whether these transactions correspond to the same or two distinct users. Thereby, RPs may have access to additional auxiliary information that might help to correlate the transaction data. Unlinkability should also hold for the same RP, i.e., a user authenticating twice or more often to the same RP. Unless the user has established a unique pseudonym with an RP, which clearly would break unlinkability trivially and intentionally.
@@
The IdP should not learn any information to which RPs a user authenticates. So even if the IdP observes all actions at the RPs, it should not be able to learn such information. This property is sometimes called unobservability.
@@
What does this mean for anonymity? We stress that unlinkability is a technical term and needs to be rigorously formalized to capture the desired intuition. For instance, it is clear that unlinkability always needs to be considered along with the revealed attributes and it can only be guaranteed among a set of users with the same attributes, i.e., the respective anonymity set which hides the identity of the user. The degree of anonymity provided in such a setting always depends on the size of the anonymity set, i.e., the set of users that could reveal the same attributes. Consequently, the more unique the combination of revealed attributes the smaller the degree of anonymity, although technically unlinkability can still be satisfied if there are at least two potential users.
@@ 3.2 Discussion of Privacy in Traditional Approaches
Signature-based authentication. For a signature-based approach, just like in password-based authentication, the (required) user attributes are just stored with each RP. Moreover, the certificate, which is required for authentication to check the authenticity of $\mathsf{pk}$ always reveals all its encoded attributes.
Single-sign-on. For the single-sign-on approach, as it is realized by OpenID Connect, we can observe the following. The IdP knows all the required attributes of the user, and when authenticating at an RP the IdP may only reveal attributes (via the ID token) that are required. But this is clearly outside the control of the user and so selective disclosure is not strictly enforced. When it comes to unlinkability, different RPs can typically track a user as the ID token contains some unique identifier of the user. The IdP clearly can link all interactions as it is the one issuing the ID token. Moreover, any collaboration of IdP and RPs can trivially link actions of one user together.
Beyond the currently deployed approaches. While not (yet) practically deployed, we however want to note that there are research works, e.g., by Lehmann et al. [DBLP:journals/popets/KroschewskiL23, cryptoeprint:2024/1124], that investigate how approaches like OpenID Connect could work with an increased privacy protection. Also, Yeoh et al. [DBLP:conf/uss/YeohKHKH23] present an approach to increase privacy in the FIDO2 protocol. We will not further discuss these approaches here but note that they use techniques, in particular anonymous credentials, that we are going to discuss subsequently.

### v2

@@ Abstract
With the increasing use of online services, the protection of the privacy of users becomes more and more important. This is particularly critical as authentication and authorization, as realized on the Internet nowadays, typically rely on centralized identity management solutions. Although those are very convenient from a user’s perspective, they are quite intrusive from a privacy perspective and are currently far from implementing the concept of data minimization. Fortunately, cryptography offers exciting primitives such as zero-knowledge proofs and advanced signature schemes to realize various forms of so-called anonymous credentials. Such primitives enable online authentication and authorization with a high level of built-in privacy protection (what we call privacy-preserving authentication). Though these primitives have already been researched for various decades and are well understood in the research community, unfortunately, they lack widespread adoption. In this paper, we look at the problems, what cryptography can do, some deployment examples, and barriers to widespread adoption, which we discuss using the example of the EU Digital Identity Wallet (EUDIW) and the recent discussion and feedback from cryptography experts around this topic. We also briefly comment on the transition to post-quantum cryptography.
@@ 1 Introduction
Authentication and authorization are two closely coupled and important tasks that are frequently required when users want to access resources in the digital realm, e.g., services on the Internet. Loosely speaking, authentication is the task of confirming whether a user is really who they pretend to be, e.g., by demonstrating the knowledge of some secret such as a password or a secret signing key associated with the user. Today, one increasingly uses a combination of more than one factor (multi-factor authentication), like the knowledge of a password and the possession of Subscriber Identity Module (SIM) card. Authorization is the task of giving users permissions to access certain resources, i.e., deciding whether a certain authenticated user has sufficient permissions to perform some action. Henceforth, we will just talk about authentication and note that typically authorization can be performed based on information associated with the identity after authentication and without any further interaction with the user. A typical example is the use of attributes associated with users, e.g., any user with attribute role=Student might be granted access to a certain service.
@@
A convenient way to think about a digital identity is that of a set of attributes that describe a person. Now, in a typical scenario a person’s identity comprises several not necessarily disjoint partial identities (cf. [anonterminology]), where each of them represents a person in a specific context or role and can be thought of as a certain subset of attributes. For instance, every person has a set of attributes related to health (e.g., health status, medical history), which they might not want to include in their partial identity for “work”. The complete identity of a person is represented by the union of all the attributes of all the partial identities of this person. Since some of these partial identities include sensitive information, e.g., the health data mentioned above, from a privacy perspective it is desirable to have full control over which attributes are revealed in a certain context. This is particularly relevant in the digital realm where it is easy to collect such information, connect it to other information sources and use it to build up ever increasing profiles of individuals.
@@
Limiting the amount of information that is provided as much as possible is known as data minimization. It is a concept of growing importance in a globally connected digitized world and a fundamental principle behind many data protection regulations such as the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States. This helps to preserve privacy and can help to ensure as much unlinkability as possible, making it harder to easily connect all the partial identities of a person. One issue that is important to stress is that for many services on the Internet, concretely identifying information such as name or DateOfBirth of a user might not be required in the process of authentication, e.g., it can be sufficient for an individual to demonstrate that they are old enough to consume a certain service. What, however, can be required in various cases such as governmental applications, is that actions of the same partial identity of a person can be linked over different sessions. This can be achieved via so-called pseudonyms, which represent identifiers other than the person’s real name (or other uniquely identifying information) that are independent of the person and provide a sufficient degree of anonymity. [^2] We stress that pseudonymity and anonymity are not used in the strict legal sense but rather in a technical sense relevant to the paper’s context.
@@
After having introduced some basic concepts, our goal is to look into how authentication on the Internet typically looks today and how it could look when putting privacy as the main design goal. We will discuss issues with these existing mechanism and present a concept from the cryptographic literature called anonymous credentials. This is an important tool to realize privacy-preserving authentication first envisioned by Chaum [DBLP:journals/cacm/Chaum85] and later realized by Camenisch and Lysyanskaya [DBLP:conf/eurocrypt/CamenischL01] as well as Brands [BrandsBook] in the early 2000s. After around 25 years, there is a large body of research on anonymous credentials and related primitives (cf. [DBLP:conf/secsr/KakviMPQ23, cryptoeprint:2023/1039]). Unfortunately, although this concept is well known in the research community, and over the recent years increasingly attracting practitioners, it still lacks a widespread deployment. Besides providing a high-level discussion of the theory of anonymous credentials, this paper will also look at the practice of anonymous authentication and potential issues that are hindering a widespread adoption. As a supporting use case, we take the European Digital Identity Wallet (EUDIW), its architectural reference framework (ARF) and the recent feedback given by cryptographic experts [cryptoeudi], which highlights obstacles on the way to deploying such a “new” technology into an existing (legacy) infrastructure.
@@
Outline of this paper. In Section 2 we review traditional authentication on the Internet. Then, in Section 3 we introduce relevant privacy properties and discuss privacy of traditional authentication approaches. In Section 4 we introduce the concept of anonymous credentials from a conceptual (theoretical) perspective. Then, in Section 5 we discuss anonymous credentials from a practical perspective and in particular implementations, standardization efforts and real-world use cases. In Section 6 we take a look at the European Digital Identity Wallet (EUDIW), its planned realization and related privacy issues. Then in Section 7 we look at barriers and issues to be considered when aiming at deploying anonymous credentials such as integrating them into the EUDIW. Finally, in Section 8 we conclude this paper and give an outlook on future aspects related to deploying anonymous credentials.
@@ 2 Traditional Authentication and Identity-Management
The very traditional way of authenticating users on the Internet is that every service realizes its own authentication mechanism, meaning that a user establishes a partial identity with every single service (typically pseudonyms or more likely email addresses). These partial identities are though typically not distinct, as users tend to re-use their email addresses either as an explicit pseudonym or as an attribute given to the service provider. The latter is usually required for functionality reasons, i.e., account recovery.
@@ 2.1 Password-based Authentication
The most common traditional authentication mechanism is via user-chosen passwords associated with user accounts. Since strong passwords are hard to remember, it is well known that users tend to choose too weak passwords [DBLP:conf/sp/Bonneau12] as well as reuse them among different services [DBLP:conf/soups/WashRBW16]. While there are password-management tools that help to overcome these issues, research shows that the use of such tools significantly lags behind [DBLP:conf/soups/ZhangPBC19] and due to various reasons does not see widespread use. Despite all these issues, due to its simplicity and easy deployment, it is still the most common authentication mechanism.
@@ 2.2 Signature-based Authentication
From a security perspective, two-factor authentication that relies on strong cryptography is more desirable and has been increasingly promoted. A popular choice is the use of the FIDO2333https://fidoalliance.org/fido2/ de-facto standard from the Fast IDentity Online (FIDO) alliance. Here one either uses a dedicated hardware token or a pure software component (typically called Passkey [^4] https://fidoalliance.org/passkeys/) as the second factor. The latter can be easily used with any device like a smartphone and thus enhances the usability.
@@
The technical protocol details are out of the scope of this paper, but the relevant part is that it uses public-key cryptography and in particular digital signatures. We recall that in the context of a digital signature scheme every user can generate a pair $(\mathsf{sk},\mathsf{pk})$ consisting of a secret (signing) key $\mathsf{sk}$ and a public (verification) key $\mathsf{pk}$, where $\mathsf{pk}$ is made public. A user knowing $\mathsf{sk}$ can produce a signature $\sigma\leftarrow\mathsf{Sign}(\mathsf{sk},m)$ for any message $m$, and using $\mathsf{pk}$ and, given $(m,\sigma)$, anyone can check the validity of the signature using $\{0,1\}\leftarrow\mathsf{Verify}(\mathsf{pk},m,\sigma)$, where $1$ indicates a valid signature. A valid signature gives the guarantees that $m$ has really been signed by the holder of the secret key corresponding to $\mathsf{pk}$ and has not been altered in any way. Security requires that a signature scheme is existentially unforgeable under a chosen message attack (i.e., provides EUF-CMA security). This means that when only having $\mathsf{pk}$ and access to signatures for arbitrarily and adaptively chosen messages, it is not possible to produce a valid signature $\sigma$ for some message $m^{\prime}$ that has not been explicitly signed by the signer, i.e., for which no signature was requested.
@@
In conventional web authentication, $\mathsf{pk}$ is typically just stored together with the identity and thereby establishing the binding between the individual and the key. More generally, this binding between $\mathsf{pk}$ and an identity is achieved via explicit certification. The most common way is to rely on a public-key infrastructure (PKI). This means that there is some trusted certification authority (CA) that signs $\mathsf{pk}$ together with a set of attributes of the user, resulting in a certificate that can be verified by anyone trusting the public key of the CA. This certificate is typically used for signature-based authentication when using an official governmental electronic identity (eID), as largely deployed within the European Union, where some governmental body acts as a CA.
2.3 Single Sign-On
@@
A widespread concept used nowadays is that of single sign-on (SSO), a concept with a rich history and prominent schemes such as Microsoft Passport (cf. [sso-pashalidis] for a comprehensive treatment of that topic). It makes account management simpler as it requires users to authenticate only at a single centralized entity (typically a large technology company such as Google, Microsoft, or Meta) acting as an identity-provider (IdP) for the user. Then one employs an authentication and authorization layer such as OpenID Connect (OIDC), which enables authentication at third-party services (so-called relying parties or RPs) via the IdP. We illustrate the process in Figure 1.
@@
Figure 1: Abstract concept of single sign-on.
Essentially, whenever a user wants to authenticate at an RP (1), the user is re-directed to its IdP (2) where the user actually authenticates using e.g., password- or signature-based authentication (3). After a successful authentication, the IdP generates a token (an ID Token in case of OIDC), signed by the IdP (4), which the user can hand over to the RP (5) as a proof that the user has authenticated to the IdP.
@@ 3 Privacy Properties and Traditional Approaches
Now having discussed ways of authenticating users on the Internet, we want to discuss privacy properties that might be desirable in such mechanisms.
@@ 3.1 Privacy Properties
We will briefly summarize the privacy properties that are most important in the context of authentication at an intuitive level. We will not take into account other important security properties of authentication mechanisms which are not relevant to privacy, e.g., that it is hard to impersonate other users. Also in our further consideration we will not consider privacy on the network layer, i.e., identifiers such as IP addresses or other metadata on the network layer that might be used to identify users. There are overlay networks [DBLP:conf/sp/SyversonGR97, DBLP:conf/uss/DingledineMS04] and in particular Tor [^5] https://www.torproject.org to achieve anonymity. They can be used to hide such information and obtain anonymity on the network layer in practice. For simplicity, we just assume that using such techniques, anonymity is realized in an ideal way. Yet we stress that in practice for such networks to be effective there are many issues that need to be considered [DBLP:conf/diau/Raymond00]. Now we list the properties:
@@
Technically, unlinkability means that different actions (or transactions) of a user cannot be linked together. We will use the terms IdP, even though it might be a certificate issuer only, and RPs as a generic term for service providers. Since in such a setting there are different parties involved, we have to consider unlinkability in settings where several of these parties collaborate in trying to break the unlinkability. Below, we will use the terminology in [cryptoeudi]:
@@
If a user authenticates to different RPs, those RPs cannot determine whether these transactions correspond to the same or two distinct users. Thereby, RPs may have access to additional auxiliary information that might help to correlate the transaction data. Unlinkability should also hold for the same RP, i.e., when a user authenticates multiple times to the same RP—of course, unless the user has indeed established a unique pseudonym that would intentionally break unlinkability.
@@
The IdP should not learn any information about which RPs a user authenticates to. So even if the IdP observes all actions at the RPs, it should not be able to learn such information. This property is sometimes called unobservability.
@@
What does this mean for anonymity? We stress that unlinkability is a technical term and needs to be rigorously formalized to capture the desired intuition. For instance, it is clear that unlinkability always needs to be considered along with the revealed attributes and it can only be guaranteed among a set of users with the same attributes, i.e., the set in which the identity of the respective user is hidden. The degree of anonymity provided in such a setting always depends on the size of this so called anonymity set, i.e., the set of users that could reveal the same attributes. Thus, a larger anonymity set implies stronger anonymity and the more unique the combination of revealed attributes, the smaller the degree of anonymity. Technically unlinkability can thus still be satisfied if there are at least two potential users, but for a small anonymity set the degree of anonymity will be low.
@@ 3.2 Discussion of Privacy in Traditional Approaches
Signature-based authentication. For a signature-based approach, just like in password-based authentication, the (required) user attributes are just stored with each RP. Besides, in this case, the certificate — which is required for authentication to check the authenticity of the public key — always reveals all attributes encoded within it.
Single sign-on. For the single sign-on approach, as it is realized by OpenID Connect, we can observe the following. The IdP knows all the required attributes of the user, and when authenticating at an RP the IdP may only reveal attributes (via the ID token) that are required. But this is clearly outside the control of the user and so selective disclosure is not strictly enforced. When it comes to unlinkability, different RPs can typically track a user as the ID token contains some unique identifier of the user. The IdP clearly can link all interactions as it is the one issuing the ID token. Moreover, any collaboration between the IdP and RPs can trivially link actions of one user together.
Beyond the currently deployed approaches. We want to mention that Kim Cameron with his Laws of Identity [^6] https://www.identityblog.com/stories/2005/05/13/TheLawsOfIdentity.pdf and the design of Windows CardSpace (Infocard) in the early 2000s already advocated and pushed for integrating of user-controlled selective disclosure into SSO solutions. Moreover, while not (yet) practically deployed, we want to note that there are recent research works, e.g., by Lehmann et al. [DBLP:journals/popets/KroschewskiL23, cryptoeprint:2024/1124], that investigate how currently used approaches like OpenID Connect could work with an increased privacy protection. Also, Yeoh et al. [DBLP:conf/uss/YeohKHKH23] present an approach to increase privacy in the FIDO2 protocol. We will not further discuss these recent approaches here but note that they use techniques, in particular anonymous credentials, that we are going to discuss subsequently.